Write balanced symbol equations, including state symbols, for the following word equations:

i) calcium oxide + water goes to calcium hydroxide
ii) magnesium carbonate goes to magnesium oxide + carbon dioxide
iii) hydrochloric acid + sodium hydroxide goes to sodium chloride + water

b) Balance the following equations:

i) CuO(s) + HCl(aq) goes to CuCl2(aq) + H2O(l)

ii) N2(g) + H2(g) goes to NH3(g)

iii) H2SO4(aq) + NaOH(aq) goes to Na2SO4(aq) + H2O(l)

iv) NH3(g) + O2(g) goes to NO(aq) + H2O(l)
1
Expert's answer
2020-02-10T03:33:28-0500

Solution.

1.

"1.CaO(solid)+H2O(l)=Ca(OH)2(aq)"

"2.MgCO3(solid)=MgO(solid)+CO2(g.)"

"3.HCl(aq)+NaOH(aq)=NaCl(aq)+H2O(l)"

2.

"1.CuO(s) + 2HCl(aq) = CuCl2(aq)+H2O(l)"

"2.N2(g)+3H2(g)=2NH3(g)"

"3.H2SO4(aq)+2NaOH(aq)=Na2SO4(aq)+2H2O(l)"

"4.4NH3(g)+5O2(g)=4NO(aq)+6H2O(l)"

Answer:

1.

"1.CaO(solid)+H2O(l)=Ca(OH)2(aq)"

"2.MgCO3(solid)=MgO(solid)+CO2(g.)"

"3.HCl(aq)+NaOH(aq)=NaCl(aq)+H2O(l)"

2.

"1.CuO(s) + 2HCl(aq) = CuCl2(aq)+H2O(l)"

"2.N2(g)+3H2(g)=2NH3(g)"

"3.H2SO4(aq)+2NaOH(aq)=Na2SO4(aq)+2H2O(l)"

"4.4NH3(g)+5O2(g)=4NO(aq)+6H2O(l)"